OpenStudy (anonymous):

Write the equation of the line that passes through (1, 3) and (4, 4) in standard form.

OpenStudy (anonymous):

Standard form is Ax+By=C where A, B, and C are are integers (just plain positive or negative nuymbers without and fractions) Can you find slope?

OpenStudy (anonymous):

is slop 1/3?

OpenStudy (anonymous):

\[slope(m)=\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

Correct

OpenStudy (shamim):

formula is\[y-y _{1}=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }(x-x _{1})\]

OpenStudy (anonymous):

now what?

OpenStudy (anonymous):

Now take one of those points and the slope and plug it into use slope intercept form

OpenStudy (shamim):

ok

OpenStudy (anonymous):

You do this to find the y int which they call b in the formula \[y=mx+b\]

OpenStudy (shamim):

\[x _{1}=1,y _{1}=3,x _{2}=4,y _{2=4}\]

OpenStudy (anonymous):

i got 2 2/3..

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

so now its y=(1/3)x+2 2/3

OpenStudy (anonymous):

thats my answer??

OpenStudy (anonymous):

2 2/3 is C of the equation Ax+By=C and no thats not the answer yeet

OpenStudy (shamim):

ok

OpenStudy (anonymous):

i have to plug that in for Ax + By = C ?

OpenStudy (shamim):

\[y-3=\frac{ 4-1 }{ 4-3 }(x-1)\]

OpenStudy (shamim):

now solve it

OpenStudy (anonymous):

Remeber A, B, and C can't have fractions in them so after you move around the equation you get y=(1/3)x+ 2 2/3 subtract 1/3 x from both sides (-1/3)x+y= 2 2/3 multiply both sides by 3 to get rid of the fractions we get -x+3y=8 <-----

OpenStudy (anonymous):

They usually don't have the x to be negative so really x-3y= -8 is a better answer

OpenStudy (anonymous):

oh! okay i got it! thank youu!!